A quantity of a certain perfect gas is compressed from an initial state of 0.085m3,1 bar to a final state of 0.034m3,3.9 bar. The specific heat at constant volume is 0.724kn/kgK,and the specific heat at constant pressure is 1.02kj/kgK.The observed temperature rise is 146K. Calculate,(a)The specific gas constant,R (b) the mass of gas present (d)the increase of initial energy of the gas.)


1
Expert's answer
2021-05-31T06:08:51-0400

"V_1 =0.085\ m³\\
V_2= 0.034\ m³\\
P_1 = 1 \ bar\\
P_2 = 3.9\ bar\\
c_v = 0.724\ kJ/kg.K
\\∆T = 146\ K"


"nc_v∆T = nRT\ln(\dfrac{V_2}{V_1})"


"R =\dfrac{c_v}{\ln(\frac{p_1}{p_2})} = \dfrac{0.724}{\ln(\frac{3.9}{1})} = 0.532\ kJ/kg.K"